OpenStudy (anonymous):

I have two problems that I have to ask what's the formula: Solve & Check: (1) 5x + 4 =-3x -12 . Solve & Check (2)4 (3m -6)= 72 + 3 (m-8) can some one help me with the formula for the above problems. Thank you.

OpenStudy (anonymous):

5x+4=-3x-12 +3x +3x 8x+4=-12 -4 -4 8x=-16/8 x=-2

OpenStudy (anonymous):

4(3m-6)=72+3(m-8) 12m-24=72+3m-24 12m-24=48+3m -3m -3m 9m-24 =48 +24 +24 9m=72/9 m=8

OpenStudy (anonymous):

OMG Thank you so much glitterchic. You came to my rescue. I can now understand the formula to the problems. Again, TKS. lisa62

OpenStudy (anonymous):

Oh, no problem! I'm glad to help you! If you have anymore questions feel free to ask!

OpenStudy (anonymous):

yes I do have one more problem that i just came across. I hope you can help. I keep coming up with answer of 25/9 when it suppose to be 50/9 can you help with the right formula for the correct answer? Simplify: (-5/6)2 (1/2)3 = (it two of -5/6) and three of (1/2). Again tks.